please help me : this is the question
find the values of k which kx^2+8x+5=o has real roots
Thankyou

Answer by stanbon(75887)

please help me : this is the question
find the values of k which kx^2+8x+5=o has real roots
=========================
The quadratic has Real Number roots if the discriminant
is greater than or equal to zero.
=========================
d = b^2-4ac = 8^2-4*k*5
----
Solve for "k":
64 -20k >= 0
20k <= 64
k <= 3.2
==============
Cheers,
Stan H.